[Ncep.nhc.nco_contacts] TSB Morning Rounds - Wed Feb 01, 2017

TSB Admin nhc.tsbadmin at noaa.gov
Wed Feb 1 14:32:25 UTC 2017


Here is today's summary of NHC computer operations:
--- Mello

----------

TAFB:
We have a work around for rebooting the AWIPS2 lx work boxes. Forecasters  
need to first log all the way out. The next step is to reboot by clicking  
on the little circle located in the far lower right corner.

The gefs and ecens ensemble means will require coordination with the init  
team later today to get the data to ingest into GFE.
HSU:
Tuna is still down until NCO can configure the new red hat build.
Other items:

NHC RFCs:


-------------- next part --------------
An HTML attachment was scrubbed...
URL: https://www.lstsrv.ncep.noaa.gov/pipermail/ncep.nhc.nco_contacts/attachments/20170201/4828a581/attachment.html 


More information about the Ncep.nhc.nco_contacts mailing list